# On the Asymptotic Mean Square Error Optimality of Diffusion Models

### v2

@@ Appendix D Lipschitz Constant for Noise Estimation
Appendix E Connection of the DM Denoiser’s Lipschitz Constant to the SNR
The Lipschitz constant of the DM estimator $f_{{{\bm{\theta}}},1:{\hat{t}}}$ in (12) has the Lipschitz constant $L_{1:{\hat{t}}}=L_{1}L_{2:{\hat{t}}}$. Utilizing the SNR representation of the DM (10) allows to reformulate
@@ Appendix E Connection of the DM Denoiser’s Lipschitz Constant to the SNR
Interestingly, the Lipschitz constant is monotonically decreasing with decreasing SNR. Since the Lipschitz constant is related to the impact of the observation onto the estimated value, the SNR-dependency of the DM denoiser’s Lipschitz constant resembles a well-known property of the CME, i.e., the impact of the observation vanishes in low SNR since it is corrupted by a lot of noise. In the limit of a purely noisy observation, the CME yields the prior mean of the distribution and thus is constant in the observation, which is in accordance with the above result.
Appendix F Proof of Theorem 4.2
Building upon the error decomposition (18), we bound the first term by utilizing Tweedie’s formula [38] and the assumption (20) as
$\left\|\mathbb{E}^{{{\bm{\theta}}}}[{\bm{x}}_{0}|{\bm{x}}_{\hat{t}}]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{\hat{t}}]\right\|=\frac{1}{\eta^{2}}\|\nabla\log p_{{\bm{\theta}}}({\bm{y}})-\nabla\log p({\bm{y}})\|\leq\Xi/\eta^{2},$ (44)
@@
$\sum_{t=1}^{{\hat{t}}-1}L_{1:t}\sigma_{t+1}N$ $=NL_{1}\sum_{t=1}^{{\hat{t}}-1}L_{2:t}\sigma_{t+1}$ (59)
$=NL_{1}\sum_{t=1}^{{\hat{t}}-1}\frac{1-\alpha_{1}}{1-\bar{\alpha}_{t}}\sqrt{\frac{\bar{\alpha}_{t}}{\alpha_{1}}(1-\alpha_{t+1})\frac{1-\bar{\alpha}_{t}}{1-\bar{\alpha}_{t+1}}}$ (60)
@@
results in (21). The result in (22) is straightforward using (21) and [45, Theorem 2], finishing the proof.
Appendix G Proof of Theorem 4.4
@@
Appendix H Inference Steps for Constant Hyperparameters
@@
Appendix I Convergence Analysis of the Diffusion Model with Re-sampling
Proposition I.1.
@@ Appendix I Convergence Analysis of the Diffusion Model with Re-sampling
Appendix J Network Architecture
We mainly use the DM architecture from [2] with minor adaptations. Specifically, we adopt the U-net architecture [75] with wide ResNets [76]. We use time-sharing of the NN parameters with a sinusoidal position embedding of the time information [77]. The DM is trained on estimating the noise with the denoising matching term in (32) but ignoring the pre-factor, cf. [2]. We have $C_{\text{init}}=16$ initial convolutional channels, which are first upsampled with factors $(1,2,4,8)$ and $(1,2,4)$ for the 1D and 2D data, respectively, and afterward downsampled likewise to the initial resolution with residual connections between the individual up-/downsampling streams. Each up-/downsampling step consists of two ResNet blocks with batch normalization, sigmoid linear unit (SiLU) activation functions, a linear layer for the time embedding input, and a fixed kernel size of $(3\times 1)$ and $(3\times 3)$ for the 1D and 2D data, respectively. As compared to [2], we leave out attention blocks as this has not shown performance gains for the denoising task. We use a batch size of $128$ and perform a random search over the learning rate. For different timesteps, we adapt the linear schedule of $\beta_{t}$ such that approximately the same range of SNR values is covered by the DM, which has shown good results in all considered setups. The individual hyperparameters for the different timesteps and the resulting SNR range of the DM are given in Table 1.
@@
Appendix K Additional Experiments and Details
Random GMM
We choose the means ${\bm{\mu}}_{k}$ where every entry is i.i.d. $\mathcal{N}(0,1/\sqrt{N})$ and random positive definite covariances following the procedure in [78], i.e., we first generate a matrix ${\bm{S}}$ where every entry is i.i.d. $\mathcal{U}(0,1)$. Then, we compute the eigenvalue decomposition of ${\bm{S}}^{\operatorname{T}}{\bm{S}}={\bm{V}}{\bm{\Sigma}}{\bm{V}}^{\operatorname{T}}$. Finally, we construct ${\bm{C}}_{k}={\bm{V}}\operatorname{diag}({\bm{1}}+{\bm{\xi}}){\bm{V}}^{\operatorname{T}}$ for all $k=1,\dots,K$, where every entry of ${\bm{\xi}}$ is i.i.d. $\mathcal{U}(0,1)$. The weights $p(k)$ are chosen i.i.d. $\mathcal{U}(0,1)$ and afterward normalized such that $\sum_{k=1}^{K}p(k)=1$.
Refer to captionRefer to caption
@@
Pre-trained GMM
Fashion-MNIST Data. In Figure 4 and 6 (left), the analysis is performed with a pre-trained GMM based on Fashion-MNIST instead of MNIST data. Essentially, the qualitative results are very similar to those of the pre-trained GMM based on MNIST but with a trend towards a larger number of $T$ until saturation of the MSE and simultaneously a more significant decrease in the MSE over the reverse process. This may lead to the conclusion that for more complex data distributions, a higher number $T$ is necessary until convergence, and the more significant the performance gap of a practically trained DM may become in comparison to the CME, which has complete knowledge of the prior PDF. Similar to earlier results, the impact of a mismatch in the SNR information (±$\pm 3\operatorname{dB}$) is negligible, highlighting the great robustness also for more complex data distributions.
Speech Data. As an additional dataset containing natural signals, we evaluate the Librispeech audio dataset [81, (CC-BY 4.0 license)], which contains $16\operatorname{kHz}$ read English speech. To limit the number of training samples for pre-training the GMM for this denoising example, we take the “test-clean” dataset and only keep samples no longer than two seconds in duration. We perform a similar pre-processing as in [83], i.e., we transform the raw waveform to the complex-valued one-sided short-time Fourier transform (STFT) domain by using a frame length of $512$ samples, a hop size of $128$ samples and a Hann window, after which we obtain $N=256$-dimensional complex-valued data samples. As input to the DM, we stack the real and imaginary parts as separate convolution channels. The pre-trained GMM is fitted to the STFT data; also, the normalized MSE is calculated in the STFT domain, for which we add complex-valued circularly-symmetric Gaussian noise to obtain complex-valued observations at different SNR levels.
Refer to captionRefer to caption
@@
Runtime and Parameter Analysis
@@
$\text{\small SNR}=-10\operatorname{dB}$ $\text{\small SNR}=0\operatorname{dB}$ $\text{\small SNR}=10\operatorname{dB}$ $\text{\small SNR}=20\operatorname{dB}$
@@
Analysis of the Lipschitz Constants
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 7 shows the analyzed Lipschitz constants from Lemma 4.1 and Appendix D and E for the linear schedule of $\beta_{t}$ described in Appendix J. Figure 7 (top left) shows the stepwise Lipschitz constant of $f_{{{\bm{\theta}}},t}:={\bm{\mu}}_{{{\bm{\theta}}},t}$, being smaller than one for all $t=2,\dots,T$ as derived in Lemma 4.1. In contrast, the Lipschitz constant of the noise estimation NN ${\bm{\epsilon}}_{{{\bm{\theta}}},t}$, cf. Appendix D, in Figure 7 (top right) has a larger value for smaller $t$, being in accordance with the findings in [72]. Figure 7 (bottom) shows the relation between the Lipschitz constant of the proposed DM estimator $f_{{{\bm{\theta}}},1:{\hat{t}}}$ from (12) and the SNR for varying ${\hat{t}}$ as analyzed in Appendix E. Indeed, the Lipschitz constant is vanishing in the low SNR and monotonically increases for larger SNR values in agreement with the insights from Appendix E.

### v4

@@ Appendix C Lipschitz Constant for Noise Estimation
Appendix D Connection of the DM Denoiser’s Lipschitz Constant to the SNR
The DM estimator $f_{{{\bm{\theta}}},1:{\hat{t}}}$ in (13) has the Lipschitz constant $L_{1:{\hat{t}}}=L_{1}L_{2:{\hat{t}}}$. Utilizing the SNR representation of the DM (11) allows to reformulate
@@ Appendix D Connection of the DM Denoiser’s Lipschitz Constant to the SNR
Interestingly, the Lipschitz constant is monotonically decreasing with decreasing SNR. Since the Lipschitz constant is related to the impact of the observation onto the estimated value, the SNR-dependency of the DM denoiser’s Lipschitz constant resembles a well-known property of the CME, i.e., the optimization of the bias-variance trade-off.
In particular, for a low SNR value, i.e., $\eta\to\infty$, the impact of the observation onto the estimated value is vanishing since $\lim_{\eta\to\infty}\text{SNR}_{\text{DM}}({\hat{t}})=0$, cf. (12), and thus $\lim_{\eta\to\infty}L_{2:{\hat{t}}}=0$, i.e., the DM estimator (13) effectively becomes a constant, regardless of the observation. The same holds for the CME, which yields the prior mean in the asymptotically low SNR regime, indicating a vanishing variance error at the expense of a higher bias, optimizing the bias-variance trade-off of the MSE (Kay,, 1993, Ch. 10). In contrast, for a high SNR value, the observation has a high impact on the estimated value, indicated by a larger Lipschitz constant in (48). In Section J.4, these findings are experimentally validated.
Appendix E Proof of Theorem 4.2
Building upon the error decomposition (22), we bound the first term by utilizing Tweedie’s formula (Efron,, 2011) and the assumption (24) as
$\left\|\mathbb{E}^{{{\bm{\theta}}}}[{\bm{x}}_{0}|{\bm{x}}_{\hat{t}}]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{\hat{t}}]\right\|=\eta^{2}\|\nabla\log p_{{\bm{\theta}}}({\bm{y}})-\nabla\log p({\bm{y}})\|\leq\eta^{2}\Xi,$ (49)
@@
$\sum_{t=1}^{{\hat{t}}-1}L_{1:t}\sigma_{t+1}N=NL_{1}\sum_{t=1}^{{\hat{t}}-1}L_{2:t}\sigma_{t+1}$ $=NL_{1}\sum_{t=1}^{{\hat{t}}-1}\frac{1-\alpha_{1}}{1-\bar{\alpha}_{t}}\sqrt{\frac{\bar{\alpha}_{t}}{\alpha_{1}}(1-\alpha_{t+1})\frac{1-\bar{\alpha}_{t}}{1-\bar{\alpha}_{t+1}}}$ (63)
@@
results in (25). The result in (26) is straightforward using (25) and (Koller et al.,, 2022, Theorem 2), finishing the proof.
Appendix F Proof of Theorem 4.4
@@
Appendix G Inference Steps for Constant Hyperparameters
@@
Appendix H Convergence Analysis of the Diffusion Model with Re-sampling
Proposition H.1.
@@ Appendix H Convergence Analysis of the Diffusion Model with Re-sampling
Appendix I Network Architecture
We mainly use the DM architecture from (Ho et al.,, 2020) with minor adaptations. Specifically, we adopt the U-net architecture (Ronneberger et al.,, 2015) with wide ResNets (Zagoruyko and Komodakis,, 2017). We use time-sharing of the NN parameters with a sinusoidal position embedding of the time information (Vaswani et al.,, 2017). The DM is trained on estimating the noise with the denoising matching term in (37) but ignoring the pre-factor, cf. Ho et al., (2020). We have $C_{\text{init}}=16$ initial convolutional channels, which are first upsampled with factors $(1,2,4,8)$ and $(1,2,4)$ for the 1D and 2D data, respectively, and afterward downsampled likewise to the initial resolution with residual connections between the individual up-/downsampling streams. Each up-/downsampling step consists of two ResNet blocks with batch normalization, sigmoid linear unit (SiLU) activation functions, a linear layer for the time embedding input, and a fixed kernel size of $(3\times 1)$ and $(3\times 3)$ for the 1D and 2D data, respectively. As compared to Ho et al., (2020), we leave out attention blocks as this has not shown performance gains for the denoising task. We use a batch size of $128$ and perform a random search over the learning rate. For different timesteps, we adapt the linear schedule of $\beta_{t}$ such that approximately the same range of SNR values is covered by the DM, which has shown good results in all considered setups. The individual hyperparameters for the different timesteps and the resulting SNR range of the DM are given in Table 1.
@@ Appendix I Network Architecture
Appendix J Additional Experiments and Details
J.1 Random GMM
We choose the means ${\bm{\mu}}_{k}$ where every entry is i.i.d. $\mathcal{N}(0,1/\sqrt{N})$ and random positive definite covariances following the procedure of Pedregosa et al., (2011), i.e., we first generate a matrix ${\bm{S}}$ where every entry is i.i.d. $\mathcal{U}(0,1)$. Then, we compute the eigenvalue decomposition of ${\bm{S}}^{\operatorname{T}}{\bm{S}}={\bm{V}}{\bm{\Sigma}}{\bm{V}}^{\operatorname{T}}$. Finally, we construct ${\bm{C}}_{k}={\bm{V}}\operatorname{diag}({\bm{1}}+{\bm{\xi}}){\bm{V}}^{\operatorname{T}}$ for all $k=1,\dots,K$, where every entry of ${\bm{\xi}}$ is i.i.d. $\mathcal{U}(0,1)$. The weights $p(k)$ are chosen i.i.d. $\mathcal{U}(0,1)$ and afterward normalized such that $\sum_{k=1}^{K}p(k)=1$.
@@
J.2 Pre-trained GMM
Fashion-MNIST Data. In Figure 4 and 6 (left), the analysis is performed with a pre-trained GMM based on Fashion-MNIST instead of MNIST data. Essentially, the qualitative results are very similar to those of the pre-trained GMM based on MNIST but with a trend towards a larger number of $T$ until saturation of the MSE and simultaneously a more significant decrease in the MSE over the reverse process. This may lead to the conclusion that for more complex data distributions, a higher number $T$ is necessary until convergence, and the more significant the performance gap of a practically trained DM may become in comparison to the CME, which has complete knowledge of the prior distribution. Similar to earlier results, the impact of a mismatch in the SNR information (±$\pm 3\operatorname{dB}$) is negligible, highlighting the great robustness also for more complex data distributions.
Speech Data. As an additional dataset containing natural signals, we evaluate the Librispeech audio dataset (Panayotov et al.,, 2015), which contains $16\operatorname{kHz}$ read English speech. To limit the number of training samples for pre-training the GMM for this denoising example, we take the “test-clean” dataset and only keep samples no longer than two seconds in duration. We perform a similar pre-processing as Welker et al., (2022), i.e., we transform the raw waveform to the complex-valued one-sided short-time Fourier transform (STFT) domain by using a frame length of $512$ samples, a hop size of $128$ samples and a Hann window, after which we obtain $N=256$-dimensional complex-valued data samples. As input to the DM, we stack the real and imaginary parts as separate convolution channels. The pre-trained GMM is fitted to the STFT data; also, the normalized MSE is calculated in the STFT domain, for which we add complex-valued circularly-symmetric Gaussian noise to obtain complex-valued observations at different SNR levels.
@@
J.3 Runtime and Parameter Analysis
@@
$\text{SNR}=-10\operatorname{dB}$ $\text{SNR}=0\operatorname{dB}$ $\text{SNR}=10\operatorname{dB}$ $\text{SNR}=20\operatorname{dB}$
@@
J.4 Analysis of the Lipschitz Constants
@@
Figure 7 shows the analyzed Lipschitz constants from Lemma 4.1 and Appendix C and D for the linear schedule of $\beta_{t}$ described in Appendix I. Figure 7 (top left) shows the stepwise Lipschitz constant of $f_{{{\bm{\theta}}},t}:={\bm{\mu}}_{{{\bm{\theta}}},t}$, being smaller than one for all $t=2,\dots,T$ as derived in Lemma 4.1. In contrast, the Lipschitz constant of the noise estimation NN ${\bm{\epsilon}}_{{{\bm{\theta}}},t}$, cf. Appendix C, in Figure 7 (top right) has a larger value for smaller $t$, being in accordance with the findings of Yang et al., (2024). Figure 7 (bottom) shows the relation between the Lipschitz constant of the proposed DM estimator $f_{{{\bm{\theta}}},1:{\hat{t}}}$ from (13) and the SNR for varying ${\hat{t}}$ as analyzed in Appendix D. Indeed, the Lipschitz constant is vanishing in the low SNR and monotonically increases for larger SNR values in agreement with the insights from Appendix D.
J.5 Qualitative Results
In this section, we provide qualitative analysis of the proposed DM-based denoiser for the MNIST and Fashion-MNIST datasets, illustrating the effect of the perception-distortion trade-off (Blau and Michaeli,, 2018) and further validating our proposed denoising approach. We note that the noise realization is not uniform across all denoisers, leading to slightly different initial points for the denoising task.
J.5.1 Datasets Based on Pre-trained GMM
We first qualitatively analyze the denoising performance on datasets sampled from a pre-trained GMM, see Section 5 and Section J.2. For each dataset, we compare the performance of our proposed DM-based denoiser with the ground-truth CME, cf. (33), and the re-sampling-based approach for multiple randomly chosen samples from the test dataset across different SNR values, i.e., different noise levels in the observations. Please note that, since the “original” images are sampled from the pre-trained GMM rather than from the original MNIST and Fashion-MNIST datasets, they may contain certain artifacts arising from the mismatch between the GMM distribution and the true bounded density of images. However, the advantage of using these data lies in the ability to evaluate the ground-truth CME, thereby enabling a clearer visualization of the perception-distortion trade-off, which is a central aspect of this experiment. Later in Section J.5.2, we also evaluate the original datasets without the ground-truth CME baseline.
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CME
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CME
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Qualitative denoising performance for different SNR levels for the the pre-trained GMM based on MNIST with $K=128$.
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CME
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CME
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Qualitative denoising performance for different SNR levels for the the pre-trained GMM based on Fashion-MNIST with $K=128$.
The qualitative analyses in Figure 8 and Figure 9 for the pre-trained GMM based on the MNIST and Fashion-MNIST datasets, respectively, show that the CME and our DM-based denoiser perform very similarly, despite differences in noise realizations for the observations. This further validates the theoretical analysis on a qualitative basis. Moreover, compared to DM-based denoising using re-sampling, the perception-distortion trade-off is clearly evident, particularly in the low SNR (high noise) regime. Specifically, the CME and our DM-based denoiser produce point estimates that appear more “blurry,” resembling a weighted average over different images—consistent with the definition of the CME. In contrast, the re-sampling-based DM denoiser generates more perceptually appealing images, even in the low SNR (high noise) regime. However, this approach may yield completely different prior samples from the original data point, ultimately leading to a higher MSE (i.e., a higher distortion metric). We refer to the discussion in Section 4.5.
J.5.2 Original Datasets
In Figure 10 and Figure 11, the underlying DM is trained directly on the original MNIST and Fashion-MNIST datasets, respectively, without relying on a pre-trained GMM as the ground-truth distribution. Consequently, no ground-truth CME baseline is available. The results further support the previous discussion on the perception-distortion trade-off and confirm the applicability of our proposed DM-based denoiser in the natural image domain. Similar to the findings in Section J.5.1, the difference between our deterministic DM-based denoiser and the re-sampling-based approach becomes more visually pronounced in the low SNR (high noise) regime. While our approach optimizes the distortion metric (i.e., minimizes the MSE), resulting in potentially more “blurry” but MSE-optimal reconstructions, the re-sampling-based method aims to generate high-quality perceptual features, which may not exist in the original sample. This ultimately leads to higher distortion (i.e., higher MSE).
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative denoising performance for different SNR levels based on the original MNIST dataset.
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SNR $-20\operatorname{dB}$ $-10\operatorname{dB}$ $0\operatorname{dB}$ $10\operatorname{dB}$ $20\operatorname{dB}$
Original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DM (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Re-samp
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Qualitative denoising performance for different SNR levels based on the original Fashion-MNIST dataset.